# Polychromic Objectives for Reinforcement Learning

### v1

@@ 6 Related Work
Policy gradient methods  [Sutton et al., 1999; Sutton & Barto, 2018; Kakade, 2001] are a widely used family of algorithms for online reinforcement learning. Numerous variants have been developed to reduce variance and improve sample efficiency  [Bhatnagar et al., 2009; Degris et al., 2013; Silver et al., 2014; Lillicrap et al., 2016; Wang et al., 2017; Schulman et al., 2017a, b; Roux et al., 2025]. More recently, in the context of LLM fine-tuning, several works have replaced the learned critic with empirical estimates  [DeepSeek-AI et al., 2025; Yu et al., 2025; Zheng et al., 2025; Liu et al., 2025]. The availability of test-time compute has motivated training objectives aligned with inference-time metrics  [Tang et al., 2025; Chow et al., 2024].
RL fine-tuning often suffers from entropy collapse, where policies concentrate on a few high-reward behaviors and lose coverage of the pretrained distribution [Cui et al., 2025; Wu et al., 2025; Yue et al., 2025]. Entropy bonuses [Haarnoja et al., 2017, 2018; Schulman et al., 2017b; Seo et al., 2021; Islam et al., 2019] mitigate this locally but may struggle to induce semantic or trajectory-level exploration. In contrast, UCB-style bonus  [He et al., 2025; Lanchantin et al., 2025] can be very effective in encouraging exploration especially earlier on during RLFT; however, the objective does not explicitly encourage preserving this exploratory behavior or diversity throughout training as the UCB bonus decays over time. Covariance controls [Cui et al., 2025] can slow down entropy collapse but it is not clear that these methods necessarily encourage the policy to explore diverse generations. Our analysis builds on top of the insights Cui et al. [2025] propose and we show that our method also exercises covariance controls on homogeneous behaviors. Intrinsic curiosity [Pathak et al., 2017] also encourage exploration but it is not clear that these methods either scale to high-dimensional spaces or solve the local versus semantic exploration problem. Other approaches, more similar to our work, explicitly reward diversity, e.g., diversity-weighted objectives [Li et al., 2025] or batch-level exploration rewards [Song et al., 2025]. The crucial distinction our work makes is the set reinforcement learning framework that enables learning a set of behaviors wherein some trajectories receive positive gradient updates for exploration despite not contributing rewards to the set.
@@ 8 Acknowledgments
We would like to thank Amber Xie, Joey Hejna, Suvir Mirchandani, Hung-Chieh Fang and Andy Tang for feedback on an earlier version of the paper. We thank Anikait Singh and Ayush Chakravarthy for helpful comments on the experimental evaluations. We thank Yoonho Lee for several valuable comments on formalizing diversity. We thank Hamidur Rahman and Pervin Akhter for helpful guidance on the formalization of the problem and feedback on the project in general. We thank Zabed Hasan and Fazilat Afreen for their insightful feedback on experimental evaluations. This work is supported by an NSF CAREER award, NSF award # 1941722, ONR YIP, DARPA YFA award # W911NF2210214 and Schmidt Sciences.
@@ Appendix A Implementation Details
For BabyAI tasks, the policy conditions on the grid image, the agent’s direction embedding, and the mission text (encoded by a GRU); the action space is the standard BabyAI/MiniGrid discrete set left, right, forward, pickup, drop, toggle, done. We provide example configurations for each task in fig. 5. We train a CNN–GRU policy that outputs action logits. In MiniGrid–FourRooms, the action space is identical, but the observation excludes a mission (as the goal specification is fixed), so we use a compact MLP that produces action logits conditioned on a flattened image observation. During pretraining, we use an 80/20 train–test split of expert demonstrations for each task, except for Synthseq and BossLevel, where we found that the full dataset was required to obtain a reasonably strong base policy. We used the dataset from Younis et al. [2024]. We pretrain by minimizing the cross-entropy loss with an entropy regularizer.
@@
In Triangle Discovery, an agent must output a sequence of edge tokens that form a valid triangle in an unobserved undirected graph. The input sequence comprises a graph index plus a prefix prompt and the agent’s past outputs. We set the maximum sequence length to be $11$. The action space is a discrete vocabulary of size $1017$. We pretrain a decoder-only Transformer with masked cross-entropy. The pretraining dataset, from Nagarajan et al. [2025], contains 15,00015{,}000 samples per graph, with the same being a triangle with probability $\tfrac{1}{3}$ and an edge with probability $\tfrac{2}{3}$ edges. Each graph has $999$ nodes. For RLFT, we fine-tune on 3 fixed graphs. The reward is sparse: +1+1 for a valid triangle, 0 otherwise.
@@ A.1.1 Vine Sampling
In this section, we describe the vine sampling scheme  [Schulman et al., 2017a] we used for polychromic PPO . In vine sampling, we first generate a number of trajectories. Then, we select a subset of the states visited, denoted by $\{s_{1},\ldots,s_{p}\}$ - this is called the rollout set. For each state $s_{i}$ in the rollout set (we call this a rollout state), we generate a set of trajectories $\tau_{1:N}\sim\pi_{\beta}(\cdot\mid s_{i})$. To avoid notational overload and to make sample accounting explicit, we distinguish:
@@
In this paper, we used the third criterion. Note that, in this scheme, we generate, in total, $N+N^{2}(p-1)$ trajectories. Since we are allowed to sample at most $B$ trajectories, we must select $N$ and $p$ such that $N>n$ and $N+N^{2}(p-1)\leq B$. Across all environments, we set a trajectory budget $B=136$ for all methods. As such, we use sets of size $n=4$ for set RL, $N=8$ vines at each rollout state, and $p=2$ rollout states per trajectory. Note that at each of these rollout states, given we generate $N$ rollouts, we can find $\binom{N}{n}$ sets for our set RL algorithm. We set the number of sets, $M$, be 4. For our chosen hyperparameters, this was sufficiently large number of sets from each rollout state allowing us to use several sets to compute the baseline.
@@ A.1.2 Other Implementation Details and Pseudocode
3:  Collect on-policy data using fractal sampling:
@@ Appendix C Proofs
Now we prove Lemma 3.2. Our proof follows the same procedure as in the proof of the standard performance difference lemma in Kakade & Langford [2002]. We first prove the following lemma:
@@
We follow the set-up in Cui et al. [2025]. We parameterize the policy as a softmax distribution:
@@
Proof. We have the following first order approximation of $\Delta$ [Cui et al., 2025]:
@@
C.3 Proof of Proposition 5.3
@@
$\geq P_{n}(1-p)\cdot\frac{q}{n}$
$=\frac{qp^{n}(1-p)}{n}$

### v2

@@ 6 Related Work
Policy gradient methods  [34, 33, 14] are a widely used family of algorithms for online reinforcement learning. Numerous variants have been developed to reduce variance and improve sample efficiency  [2, 7, 19, 37, 27, 29, 26]. More recently, in the context of LLM fine-tuning, several works have replaced the learned critic with empirical estimates  [6, 41, 45, 20]. The availability of test-time compute has motivated training objectives aligned with inference-time metrics  [36, 4].
RL fine-tuning often suffers from entropy collapse, where policies concentrate on a few high-reward behaviors and lose coverage of the pretrained distribution [5, 39, 42]. Entropy bonuses [9, 10, 29, 30, 12] mitigate this locally but may struggle to induce semantic or trajectory-level exploration. In contrast, UCB-style bonus  [11, 17] can be very effective in encouraging exploration especially earlier on during RLFT; however, the objective does not explicitly encourage preserving this exploratory behavior or diversity throughout training as the UCB bonus decays over time. Covariance controls [5] can slow down entropy collapse but it is not clear that these methods necessarily encourage the policy to explore diverse generations. Our analysis builds on top of the insights [5] propose and we show that our method also exercises covariance controls on homogeneous behaviors. Intrinsic curiosity [24] also encourage exploration but it is not clear that these methods either scale to high-dimensional spaces or solve the local versus semantic exploration problem. Other approaches, more similar to our work, explicitly reward diversity, e.g., diversity-weighted objectives [18], outcome-based UCB bonus [32] or train explicitly for exploration using synthetic data  [35]. The crucial distinction our work makes is the set reinforcement learning framework that enables learning a set of behaviors wherein some trajectories receive positive gradient updates for exploration despite not contributing rewards to the set.
@@ 8 Acknowledgments
We would like to thank Amber Xie, Joey Hejna, Suvir Mirchandani, Hung-Chieh Fang and Andy Tang for feedback on an earlier version of the paper. We thank Anikait Singh and Ayush Chakravarthy for helpful comments on the experimental evaluations. We thank Yoonho Lee for several valuable comments on formalizing diversity. We thank Hamidur Rahman and Pervin Akhter for helpful guidance on the formalization of the problem and feedback on the project in general. We thank Zabed Hasan and Fazilat Afreen for their insightful feedback on experimental evaluations. This work is supported by an NSF CAREER award, NSF award # 1941722, ONR YIP, DARPA YFA award # W911NF2210214, and Schmidt Sciences.
@@ Appendix A Implementation Details
For BabyAI tasks, the policy conditions on the grid image, the agent’s direction embedding, and the mission text (encoded by a GRU); the action space is the standard BabyAI/MiniGrid discrete set left, right, forward, pickup, drop, toggle, done. We provide example configurations for each task in Fig. 5. We train a CNN–GRU policy that outputs action logits. In MiniGrid–FourRooms, the action space is identical, but the observation excludes a mission (as the goal specification is fixed), so we use a compact MLP that produces action logits conditioned on a flattened image observation. During pretraining, we use an 80/20 train–test split of expert demonstrations for each task, except for Synthseq and BossLevel, where we found that the full dataset was required to obtain a reasonably strong base policy. We used the dataset from [40]. We pretrain by minimizing the cross-entropy loss with an entropy regularizer.
@@
In Triangle Discovery, an agent must output a sequence of edge tokens that form a valid triangle in an unobserved undirected graph. The input sequence comprises a graph index plus a prefix prompt and the agent’s past outputs. We set the maximum sequence length to be $11$. The action space is a discrete vocabulary of size $1017$. We pretrain a decoder-only Transformer with masked cross-entropy. The pretraining dataset, from [21], contains 15,00015{,}000 samples per graph, with the same being a triangle with probability $\tfrac{1}{3}$ and an edge with probability $\tfrac{2}{3}$ edges. Each graph has $999$ nodes. For RLFT, we fine-tune on 3 fixed graphs. The reward is sparse: +1+1 for a valid triangle, 0 otherwise.
@@ A.1.1 Vine Sampling
In this section, we describe the vine sampling scheme  [27] we used for polychromic PPO . In vine sampling, we first generate a number of trajectories. Then, we select a subset of the states visited, denoted by $\{s_{1},\ldots,s_{p}\}$ - this is called the rollout set. For each state $s_{i}$ in the rollout set (we call this a rollout state), we generate a set of trajectories $\tau_{1:N}\sim\pi_{\beta}(\cdot\mid s_{i})$. To avoid notational overload and to make sample accounting explicit, we distinguish:
@@
In this paper, we used the third criterion. Note that, in this scheme, we generate, in total, $N+N^{2}(p-1)$ trajectories. Since we are allowed to sample at most $B$ trajectories, we must select $N$ and $p$ such that $N>n$ and $N+N^{2}(p-1)\leq B$. Across all environments, we set a trajectory budget $B=136$ for all methods. As such, we use sets of size $n=4$ for set RL, $N=8$ vines at each rollout state, and $p=2$ rollout states per trajectory. Note that at each of these rollout states, given we generate $N$ rollouts, we can find $\binom{N}{n}$ sets for our set RL algorithm. In fact, this combinatorial construction of sets allows us to compute a baseline that is unbiased. In our experiments, we set the number of sets, $M$, to be 4. For our chosen hyperparameters, this was sufficiently large number of sets from each rollout state allowing us to use several sets to compute the baseline.
@@ A.1.2 Other Implementation Details and Pseudocode
3:  Collect on-policy data using vine sampling:
@@ Appendix C Proofs
Now we prove Lemma 3.2. Our proof follows the same procedure as in the proof of the standard performance difference lemma in [13]. We first prove the following lemma:
@@
We follow the set-up in [5]. We parameterize the policy as a softmax distribution:
@@
Proof. We have the following first order approximation of $\Delta$ [5]:
@@
C.3 Proof of Lemma 5.3
Lemma 5.3 (restated). Consider any set of actions $a_{1:n}\in\mathcal{A}^{n}$. The change in the log probability of sampling this set of actions after one policy update using set RL can be written as the following first-order approximation:
$\log\pi_{\theta}^{k+1}(a_{1:n}\mid s)\approx\log\pi_{\theta}^{k}(a_{1:n}\mid s)+\lambda\Lambda_{f}(a_{1:n};\pi_{\theta}^{k})-\lambda^{\prime}C(\theta^{k})$
where $C(\theta^{k})$ is a function independent of $a_{1:n}$.
Proof. This follows from using a first-order Taylor approximation of $\sum_{i=1}^{n}\log\pi_{\theta}^{k+1}(a_{i}\mid s)$ at $\sum_{i=1}^{n}\log\pi_{\theta}^{k}(a_{i}\mid s)$ and then using Lemma C.2. Proceeding this way, we get that,
$C(\theta^{k})=\mathrm{Cov}_{a_{1:n}^{\prime}\sim\pi_{\theta}^{k}(\cdot\mid s)}\left(f(s,a_{1:n}^{\prime}),\sum_{i=1}^{n}\pi_{\theta}^{k}(a_{1:n}^{\prime}\mid s)\right).$
$\square$
C.4 Proof of Proposition 5.4
@@
$\geq P_{n}(1-p)\cdot\frac{q}{n}=\frac{qp^{n}(1-p)}{n}$